# AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL

### v2

@@ IV System Overview
Figure 5: An overview of AQORA, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
  1. 1.
@@
    When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model. In Fig. 5, the system has just completed scanning table t and observed that the output cardinality is only 1. The planner extension then sends the partially executed plan and the row count (1) to the decision model. This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
  2. 2.
@@
    The decision model analyzes the input plan and returns optimization actions. As shown in Fig. 5, the decision model observes that table t produces only one output row but is joined after two larger tables. The actor therefore infers that placing t earlier in the join order could be more efficient and outputs the action lead(t).
  3. 3.
[…]
@@
4) After the query completes after $k$ iterations of 1)-3), the decision model replays the collected $k$ step optimization process to learn from past decisions and improve its optimization strategy for future queries.
We formalize this optimization process as a trajectory:
$\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$
@@
where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding cardinalities $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$.
V Decision Model
In this section, we describe the decision model in detail.
V-A The Actor-Critic Framework
@@
We adopt the actor–critic framework to fully exploit the frequent feedback signals from AQE, which promotes faster convergence and allows us to observe the critic’s reward or penalty for each action taken by the actor.
V-A1 State, Action, and Return
@@
State
Each state $s_{t}=(p_{t},c_{t})$ consists of:
  • $p_{t}$: the partial executed plan at time step $t$;
  • $c_{t}$: the true cardinalities of completed plan segments, obtained from AQE statistics.
Action Space
AQORA supports four types of actions: turning on/off CBO, manipulating join orders, setting broadcast join hints on leaf nodes to enforce broadcast joins, and performing a no-op. The notations and definitions of these actions are summarized in Tab. I:
TABLE I: Action Space of AQORA’s Decision Model
@@
cbo(1/0)(1/0) turn on/off CBO
@@
no-op do nothing
The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization. In some cases, enabling CBO produces a better starting plan that is closer to the optimal; in others, disabling it yields better results. Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method. The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front. Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
Return
Based on empirical validation, we define the return as:
$R(\tau)=\sum_{i=1}^{k}\gamma^{i-1}r_{i}-\sqrt{T_{\text{execute}}(\tau)}$
@@
where $T_{\text{execute}}$ represents the total execution time of $\tau$ (in seconds). The square root operation helps mitigate reward scale variations across different queries. If query execution fails, $R(\tau)$ is assigned a substantial negative penalty (e.g., $\sqrt{300}$). The intermediate reward $r_{i}$ is defined as the negative value of the additional number of shuffles introduced after taking action $a_{i}$, normalized by 10. This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process. The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $R(\tau)$.
V-A2 Policy and State Transition
@@
$\pi_{\theta}(a|s)=\begin{cases}P_{1},\quad&\text{if}\ a=a_{1}\ (\text{e.g.,swap}\ (1,2))\\ P_{2},\quad&\text{if}\ a=a_{2}\ (\text{e.g.,broadcast}\ (1))\\ &\vdots\\ P_{m},\quad&\text{if}\ a=a_{m}\ (\text{e.g., no-op})\end{cases}$
where $\sum_{i=1}^{m}P_{i}=1$. The probabilities $P_{i}$ are computed by the actor network’s parameters $\theta$ via softmax normalization over its output logits. This probabilistic approach enables exploring alternative optimization paths while maintaining likely good choices.
@@
The uncertainty arises from AQE’s non-deterministic triggering: the execution engine may complete a varying number of stages between optimizations. As a result, applying the same action $a$ in state $s$ can lead to different successor states $s^{\prime}$.
V-A3 State Value and Action Value
@@
V-A4 Training Algorithm
For the actor, our goal is to maximize the expectation of $q(s_{t},a_{t})$ for each step. So we use the clipped PPO loss [34] with an entropy factor to encourage explorative actions:
@@ V-A4 Training Algorithm
We observe that the policy gradient for updating the actor depends on $q$ (calculated from the critic’s state values), while the critic’s update gradient relies on $v_{\pi}$ (determined by the actor’s current policy).
As shown in Alg. 1, this bootstrapping interaction allows both components to progressively improve through repeated optimization iterations, where $\gamma$ is set to 1 for simplicity.
Algorithm 1 The training algorithm in AQORA
@@
3:Collect trajectory $(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$ using current policy $\pi_{\theta}$
@@
14:  Update $w\leftarrow w-\beta\nabla_{w}L^{\text{critic}}$
@@ V-B Model Design
When the decision model receives a partial plan $s$, it first compresses the plan tree and encodes each of its nodes into embedding vectors. The resulting vectorized tree is then processed separately by the actor and critic networks. The actor generates a probability distribution $\pi_{\theta}(s)$ over the set of valid actions, from which an action is selected using masked probabilities to filter out illegal operations. Meanwhile, the critic produces a scalar estimate $v_{\phi}(s)$, representing the state’s optimization potential.
V-B1 Tree Compressing
While nodes such as sorts and aggregates do affect query performance, we exclude them from the model’s input features to focus solely on join operations (Fig. 6(1)). This approach, though risky for other LQOs relying on full plan details, becomes viable because AQORA’s access to real-time cardinalities compensates for the lost information.
Figure 6: Tree compressing and feature encoding.
V-B2 Feature Encoding
The compressed plan tree is transformed into a vector tree, as shown in Fig. 6(2). Each node $u$ is encoded as a vector by concatenating three components:
@@ V-B2 Feature Encoding
  • Each node is assigned a type, such as a join or a broadcast query stage. AQORA encodes the node type as a one-hot vector type(u)(u). In a simplified example shown in Fig. 6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
  • $\text{table}(u)$ is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Fig. 6. It encodes the join sequence leading to node $u$, and during AQE, even leaf nodes may touch multiple tables.
[…]
This encoding strategy ensures that AQORA remains operational under data updates and schema changes. Specifically, the operator type $\text{type}(u)$ is fixed, while the intermediate cardinality $\text{card}(u)$ is collected online during execution rather than relying on precomputed statistics. Even when new tables are introduced, the encoding remains valid, with the corresponding positions taking a default value of 0.
The encoded vector tree is then processed by AQORA’s actor and critic networks, and we choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
V-B3 Masked Output
@@
for switching on/off CBO, n1n-1 possible leads (if we only consider lead a single table each time), $C_{n}^{2}$ possible swaps (each concrete action instance is treated as a distinct action, e.g., $\text{swap}(4,7)$), $n$ possible tables to broadcast, and 1 no-op. For example, when processing a partially executed plan with 3 leaf nodes, we create a binary mask as follows:
$\begin{split}M=(&\underbrace{0,0}_{\text{cbo(1), cbo(0), }},\underbrace{1,1,0,\cdots,1,0}_{\text{swap(1,2), swap(1,3), swap(2,3)}},\\ &\underbrace{1,1,0,\cdots}_{\text{lead(2), lead(3)}},\underbrace{1,1,1,0,\cdots}_{\text{broadcast(1), broadcast(2), broadcast(3)}},\underbrace{1}_{\text{no-op}})\end{split}$
However, to achieve smoother policy convergence, we implemented a curriculum learning framework that requires masking out additional actions at different training stages. Initially, the model is only allowed to choose whether to enable or disable CBO, focusing on a simple binary decision task. During the mid-training phase, we progressively lift the masking on runtime plan adjustment actions; as these decisions are based on actual runtime statistics, they are generally easier for the model to learn. In the final training stage, all restrictions are removed except for masking invalid actions, allowing the model to explore the full action space and refine its decision-making capabilities.
The final action probabilities are calculated by normalizing the masked output:
@@
VI Planner Extension
Our planner extension operates through two core mechanisms: capturing and sending the current partial plan to the decision model and applying optimization actions returned by the actor to modify the query plan. As detailed at the end of Sec. VI-B1, these mechanisms support dynamic exploration of bushy joins through iterative plan modifications at runtime.
VI-A Send the Extracted Plan
The planner extension introduces a new optimization rule into existing rule set. This rule is triggered after all default (adaptive) optimization rules have been applied. Upon activation, it first checks whether the maximum number of allowed optimization steps has been reached. If not, it extracts the current partial execution plan along with available runtime cardinalities from each completed stage. The extracted plan and cardinalities are then sent to the decision model, after which the system awaits optimization instructions.
@@ VI-A Send the Extracted Plan
VI-B Inject Plan
The optimization instructions may involve five types of operations, as shown in Tab. I. While actions such as enabling CBO and no-op operations are straightforward to handle, different strategies are required for processing join reorder and broadcast instructions.
VI-B1 Plan Transformation
Swapping instructions are denoted as swap(i,j)(i,j), indicating the exchange of positions between the $i$-th and $j$-th leaf nodes. However, if the opposite sides of the join of these two leaf nodes lack compatible join keys, executing the swap would result in a Cartesian product. To mitigate this, we design Alg. 2 to perform controlled node swapping. Leading actions follow a similar strategy.
Algorithm 2 Swapping Algorithm
1:Plan $p$ with leaf nodes $l_{1},l_{2},\cdots,l_{n}]$ and join conditions $\{c_{1},c_{2},\dots,c_{m}\}$, swap instruction swap(i,j)(i,j) or lead instruction lead(i)(i)

### v3

@@ 3.3. Spark AQE’s Limitations
Figure 3. An overview of LQRS, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
3.4. Our Solution
LQRS is tightly integrated with Spark SQL’s query execution model. It treats the completion of a query stage as a natural decision point, where partially executed plans and query stage-level statistics are exposed to the decision model. By aligning optimization actions with query stage boundaries, LQRS ensures that runtime plan adaptations are both semantically valid and practically efficient within Spark SQL’s execution framework.
Next, we introduce the main components of LQRS and its basic workflow. As shown in Figure 3, LQRS consists of two core components: a planner extension and a decision model. The figure illustrates an example where LQRS performs a single optimization action during the execution of the STACK query q1#1.
  1. (1)
@@ 3.4. Our Solution
    When Spark SQL’s AQE is triggered during execution, the planner extension extracts (①) the current query plan along with actual cardinality statistics and sends (②) it to the decision model. In Figure 3, the system has just completed scanning table t and observed the output cardinality. The planner extension then sends the partially executed plan and the row count (1) to the decision model. This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
  2. (2)
@@
    The decision model analyzes the input plan and returns (③) optimization actions. As shown in Figure 3, the decision model observes that table t produces only one output row but is joined after two larger tables. The actor therefore infers that placing t earlier in the join order could be more efficient and outputs the action lead(t).
  3. (3)
    The planner extension applies (④) these actions to the ongoing query plan and injects (⑤) the optimized plan back to the AQE engine to resume execution. As illustrated in Figure 3, after receiving the action lead(t), the planner extension attempts to reposition table t lower in the plan, allowing it to participate in the join operation earlier. In coordination with the planner extension’s reposition behavior, the AQE engine can further adapt the join strategies, transforming the original SMJ–BHJ–BHJ sequence into a more efficient BHJ–SMJ–BHJ sequence.
  4. (4)
@@
    After the query completes after $k$ iterations of ①-⑤, the decision model replays (⑥) the collected $k$ step optimization process to learn from past decisions and improve its optimization strategy for future queries. During this replay process, the critic evaluates (⑦) each action taken by the actor and provides feedback signals, which are used to guide policy improvement.
We formalize the optimization process of LQRS as an episode:
$\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{n-1},r_{k},s_{k})$
@@
where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding statistics $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$. Details will be given in the Section 4.1.1.
3.5. Our Result
As a result, compared to Spark AQE, LQRS reduces the runtime from 33.53s to 15.24s in the above example, achieving over 50% lower execution latency. In this example, Spark AQE is unable to reorder joins based on runtime observations and thus continues to execute the query using a suboptimal join order determined at pre-execution time, which both incurs higher execution cost and generates large intermediate results, preventing effective use of dynamic join selection. In contrast, LQRS captures runtime signals that reveal a better join order and proactively performs join reordering during execution, after which the AQE engine replans the remaining query with a more efficient join strategy.
4. Decision Model
4.1. The Actor-Critic Framework
@@
We adopt the actor–critic framework to fully exploit the frequent feedback signals from runtime, which promotes faster convergence and allows us to observe the critic’s reward or penalty for each action taken by the actor.
4.1.1. State, Action, and Reward
@@
State. Each state $s_{t}=(p_{t},c_{t})$ consists of:
  • $p_{t}$: the partially executed plan at time step $t$;
  • $c_{t}$: the true cardinalities and output size in bytes of completed query stages, obtained from AQE statistics.
Action Space. LQRS supports five actions: controlling the plan initialization strategy, swapping the $i$-th and $j$-th leaf node, joining the leaf nodes $t_{1},t_{2},\cdots,t_{l}$ first if possible, setting broadcast join hints on leaf nodes to enforce broadcast joins, and intentionally taking no action for now. The notations and definitions of these actions are summarized in Table 2.
Table 2. Action space of LQRS’s decision model
@@
init(1/0)(1/0) select plan initialization strategy
@@
no-op()() take no action for now
  • The init action controls the initial plan construction strategy, determining how the starting execution plan is generated before any runtime adaptation takes place. In some cases, a traditional cost-based initialization produces a plan that is already close to optimal, while in others, a rule-based initialization offers a better starting point for subsequent optimization actions. Allowing the decision model to dynamically select between different plan initialization strategies helps LQRS to escape local optima and discover better execution plans.
  • The swap action enables fine-grained join reordering by exchanging two leaf nodes in the current logical plan. This action provides local flexibility for adjusting join order and is particularly effective when small structural changes are sufficient to improve performance.
  • The lead action promotes a specific table, or a subset of tables, to be joined earlier in the execution plan. Compared to the swap action, this action offers a more direct and efficient mechanism for correcting suboptimal join orders, as many beneficial transformations can be achieved by repositioning a single critical leaf node. As a result, the lead action serves as a practical control for guiding join order decisions.
  • The broadcast action enforces a broadcast join on a selected leaf node. This action is designed to overcome the overly conservative behavior induced by static broadcast thresholds. Although forcing a broadcast join introduces some potential risks, it enables the correction of conservative join method choices.
  • The no-op action allows the decision model to intentionally defer optimization decisions and wait for additional runtime information. This action is particularly useful in early execution stages where limited feedback is available and premature interventions may be harmful. By supporting deliberate inaction, LQRS can balance exploration and stability while reducing unnecessary plan perturbations.
Together, these actions enable LQRS to perform both coarse-grained and fine-grained plan adaptations across different stages of query execution.
Reward. A key challenge of runtime plan adaptation is that modifying an execution plan after some query stages have already materialized may require redistributing intermediate data. In Spark, this typically manifests as additional shuffle operations. For example, consider a query whose initial execution plan decides to evaluate $t_{1}\bowtie t_{2}$ first, followed by a join with $t_{3}$. During execution, Spark scans $t_{1}$ and $t_{2}$ and performs a shuffle to partition their records according to the join key of $t_{1}\bowtie t_{2}$, preparing for a shuffle-based join. If a subsequent runtime decision swaps the join order to instead evaluate $t_{1}\bowtie t_{3}$ first, the existing partitioning of $t_{1}$, which was produced to satisfy the join with $t_{2}$, is no longer compatible with the new join requirement. As a result, $t_{1}$ is reshuffled to match the join key of $t_{1}\bowtie t_{3}$, allowing execution to proceed correctly under the updated plan.
Motivated by this observation, LQRS explicitly incorporates reshuffles as intermediate feedback in the reward definition, encouraging the agent to account for the operational cost of reorganizing intermediate data when making runtime decisions. Based on empirical validation, we define the total reward as:
$R(\tau)=\sum_{i=1}^{k}\gamma^{i-1}r_{i}-\sqrt{T_{\text{execute}}(\tau)},$
where $T_{\text{execute}}(\tau)$ denotes the total execution time of episode $\tau$ (in seconds). The square root operation is used to reduce reward scale variations across queries. If execution fails due to OOM errors or exceeds the time limit of 300s, $R(\tau)$ is assigned a substantial negative value (e.g., $\sqrt{300}$).
@@
The intermediate reward $r_{i}$ is defined as the negative number of additional shuffle operations introduced after taking action $a_{i}$, normalized by 10. This normalization aligns the scale of intermediate rewards with that of $\sqrt{T_{\text{execute}}(\tau)}$ under our experimental settings, ensuring that intermediate and terminal rewards are of comparable magnitude during training. This formulation naturally provides positive feedback when the agent selects a no-op action (i.e., no additional shuffle is introduced), encouraging conservative behavior in early training, while penalizing actions that trigger reshuffles and discouraging costly plan changes at later stages. The discount factor $\gamma<1$ further ensures that accumulated intermediate penalties do not dominate the overall reward.
4.1.2. Policy and State Transition
@@
$\pi_{\theta}(a|s)=\begin{cases}P_{1},\quad&\text{if}\ a=a_{1}\ (\text{e.g.,swap}(1,2))\\ P_{2},\quad&\text{if}\ a=a_{2}\ (\text{e.g.,broadcast}(1))\\ &\vdots\\ P_{m},\quad&\text{if}\ a=a_{m}\ (\text{e.g., no-op}())\end{cases}$
where $\sum_{i=1}^{m}P_{i}=1$. The probabilities $P_{i}$ are computed by the actor network with parameters $\theta$ via softmax normalization over its output logits. This probabilistic approach enables exploring alternative optimization paths while maintaining likely good choices.
@@
The uncertainty arises from Spark AQE’s non-deterministic triggering: the execution engine may complete a varying number of stages between optimizations. As a result, applying the same action $a$ in state $s$ can lead to different successor states $s^{\prime}$.
4.1.3. State Value and Action Value
@@
4.1.4. Training Algorithm
For the actor, our goal is to maximize the expectation of $q(s_{t},a_{t})$ for each step. So we use the clipped PPO loss (Schulman et al., 2017) with an entropy factor to encourage explorative actions:
@@ 4.1.4. Training Algorithm
We observe that the policy gradient for updating the actor depends on $q$ (calculated from the critic’s state values), while the critic’s update gradient relies on $v_{\pi}$ (determined by the actor’s current policy). As shown in Algorithm 1, this bootstrapping interaction allows both components to progressively improve through repeated optimization iterations, where $\gamma$ is set to 1 for simplicity.
Algorithm 1 The training algorithm in LQRS
@@
3:Collect episode $(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$ using current policy $\pi_{\theta}$
@@
14:  Update $\phi\leftarrow\phi-\beta\nabla_{\phi}L^{\text{critic}}$
@@ 4.2. Model Design
When the decision model receives a partial plan $s$, it first encodes each of its nodes into embedding vectors. The resulting vectorized tree is then processed separately by the actor and critic networks. The actor generates a probability distribution $\pi_{\theta}(s)$ over the set of valid actions. Meanwhile, the critic produces a scalar estimate $v_{\phi}(s)$, representing the state’s optimization potential.
Figure 4. Feature encoding.
4.2.1. Feature Encoding
The plan tree is transformed into a vector tree, as shown in Figure 4. Each node $u$ is encoded as a vector by concatenating three components:
@@ 4.2.1. Feature Encoding
  • type(u)(u) is a one-hot vector encoding the type of node $u$, such as a join or a query stage. In the simplified example shown in Figure 4, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0).
  • table(u)(u) is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Figure 4. It encodes the join sequence leading to node $u$, and during AQE, even leaf nodes may touch multiple tables.
  • card(u)(u) is set to $(\log(1+c))$ if $u$ corresponds to a finished query stage with observed cardinality $c$; otherwise, card(u)(u) is set to (1)(-1). The same encoding strategy is applied to other statistics, such as size in bytes.
This encoding strategy ensures that LQRS remains operational under data updates and schema changes. Specifically, the operator type $\text{type}(u)$ is fixed, while the intermediate cardinality $\text{card}(u)$ is collected online during execution rather than relying on precomputed statistics. Even when new tables are introduced, the encoding remains valid, with the corresponding positions taking a default value of 0.
The encoded vector tree is then processed by LQRS’s actor and critic networks, both of which are instantiated using a TreeCNN architecture. We refer the reader to  (Marcus et al., 2019) for a deeper investigation into tree convolution applied to query plans.
4.2.2. Masked Output
@@
for 2 init actions, n1n-1 possible lead actions(if we only consider lead a single table each time), $C_{n}^{2}$ possible swap actions (each concrete action instance is treated as a distinct action, e.g., $\text{swap}(4,7)$), $n$ possible broadcast actions, and 1 no-op action. For example, when processing a partially executed plan with 3 leaf nodes, we create a binary mask as follows:
$\begin{split}M=(&\underbrace{0,0}_{\text{init(1), init(0), }},\underbrace{1,1,0,\cdots,1,0}_{\text{swap(1,2), swap(1,3), swap(2,3)}},\\ &\underbrace{1,1,0,\cdots}_{\text{lead(2), lead(3)}},\underbrace{1,1,1,0,\cdots}_{\text{broadcast(1), broadcast(2), broadcast(3)}},\underbrace{1}_{\text{no-op}})\end{split}$
However, to achieve smoother policy convergence, we implemented a curriculum learning framework that requires masking out additional actions at different training stages. Initially, the model is only allowed to choose the plan initialization strategy, focusing on a simple binary decision task. During the mid-training phase, we progressively lift the masking on runtime plan adjustment actions; as these decisions are based on actual runtime statistics, they are generally easier for the model to learn. In the final training stage, all restrictions are removed except for masking invalid actions, allowing the model to explore the full action space and refine its decision-making capabilities. The final action probabilities are calculated by normalizing the masked output:
@@
5. Planner Extension
Our planner extension operates through two core mechanisms: extracting and sending the current partial plan to the decision model and applying optimization actions returned by the actor to modify the query plan.
The planner extension is built on top of Spark AQE. In particular, the planner extension can rely on AQE’s native support for pausing query execution at well-defined synchronization points and triggering re-optimization based on runtime statistics. In Spark AQE, a series of built-in optimization rules (including dynamic join selection) is applied in a pipelined fashion each time AQE is triggered. Our planner extension is integrated as an extension rule of Spark AQE to implement five optimization actions, without requiring the planner extension to explicitly manage execution suspension or execution state preservation.
5.1. Send the Extracted Plan
Upon triggering the planner extension rule, it first checks whether the maximum number of allowed optimization steps has been reached. If not, it extracts the current partially executed plan along with available runtime statistics from each completed query stage. The extracted plan and statistics are then sent to the decision model, after which the planner extension awaits further actions.
@@ 5.1. Send the Extracted Plan
5.2. Inject Plan
While init action and no-op action are straightforward to handle, different strategies are required for the swap action, lead action, and the broadcast action.
5.2.1. Plan Transformation
The swap action is denoted as swap(i,j)(i,j), indicating the exchange of positions between the $i$-th and $j$-th leaf nodes. However, if the opposite sides of the join of these two leaf nodes lack compatible join keys, executing the swap would result in a Cartesian product. The same issue also occurs while applying the lead action. To mitigate this, we design Algorithm  2 to perform controlled plan transformation.
Algorithm 2 The plan transformation algorithm in LQRS
1:Plan $p$ with leaf nodes $l_{1},l_{2},\cdots,l_{n}]$ and join conditions $\{c_{1},c_{2},\dots,c_{m}\}$, swap action swap(i,j)(i,j) or lead action lead(i)(i)